# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ IV-A2 DoF for 2D Structures
In this second case, ${\cal{S}}$ is a surface of size $L_{\text{x}}\times L_{\text{y}}$. Similarly to the 1D case, the current density can be represented as a 2D series expansion of Fourier basis functions. In the wavenumber domain, we have
${\widetilde{J}}(\kappa_{x},\kappa_{y})=\sum_{n_{x},n_{y}}j_{n_{x},n_{y}}\,\sqrt{L_{\text{x}}\,L_{\text{y}}}\,\text{Sinc}\left({\frac{L_{\text{x}}}{2\pi}\left(\kappa_{x}-k_{n_{x}}\right)}\right)\text{Sinc}\left({\frac{L_{\text{y}}}{2\pi}\left(\kappa_{y}-k_{n_{y}}\right)}\right)$ (24)
where $k_{n_{x}}=2\,n_{x}\,\pi/L_{\text{x}}$, and $k_{n_{y}}=2\,n_{y}\,\pi/L_{\text{y}}$ Note that only the set of indexes $(n_{x},n_{y})\in{\cal{P}}_{2}$, with ${\cal{P}}_{2}=\left\{(n_{x},n_{y}):\left(\frac{n_{x}\lambda}{L_{\text{x}}}\right)^{2}+\left(\frac{n_{y}\lambda}{L_{\text{y}}}\right)^{2}\leq 1\right\}$, satisfying $k_{n_{x}}^{2}+k_{n_{y}}^{2}\leq k_{0}^{2}$, are not evanescent, i.e., lie in the visible region. For instance, for a square surface of size $L_{\text{x}}=L_{\text{y}}=L$, it is $DoF\simeq|{\cal{P}}_{2}|\simeq\frac{\pi L^{2}}{\lambda^{2}}$ (per polarization) [15]. Interestingly, this is $\pi/4=0.79$ of the square of the DoF of a linear segment. As a consequence, by drawing a parallel as before with uniform planar arrays, we have that a $N\times N$ planar array will not be able to generate $N^{2}$ orthogonal beams (i.e., $\frac{4L^{2}}{\lambda^{2}}$), as could be erroneously suggested by intuition, but only the subset of $\frac{\pi}{4}N^{2}\approx 0.79\,N^{2}$ beams satisfying the condition of non-evanescent modes.
@@ IV-A3 DoF for 3D Structures
${\widetilde{J}}(\kappa_{x},\kappa_{y},\kappa_{z})=\sum_{n_{x},n_{y},n_{z}}j_{n_{x},n_{y},n_{z}}\sqrt{L_{\text{x}}L_{\text{y}}L_{\text{z}}}\,\text{Sinc}\left({\frac{L_{\text{x}}}{2\pi}\left(\kappa_{x}-k_{n_{x}}\right)}\right)\text{Sinc}\left({\frac{L_{\text{y}}}{2\pi}\left(\kappa_{y}-k_{n_{y}}\right)}\right)\text{Sinc}\left({\frac{L_{\text{z}}}{2\pi}\left(\kappa_{z}-k_{n_{z}}\right)}\right)$ (25)
@@
Figure 6: Filtering operation of the Green propagation operator in the wavenumber domain for $L/\lambda=6$. The sphere corresponding to the set ${\cal{E}}$ is depicted in blue, while the grey points represent the 3D Sinc functions in (25). The red points instead denote the subset of 3D Sinc functions in (25) having significant intersection with the sphere, hence generating EM waves.
The derivation of the number of DoF for a 3D structure is more involved. The typical approach is to consider the smallest sphere enclosing the space ${\cal{S}}$, define a complete basis set given by the spherical harmonic functions on its surface, and identify the $N$ dominant ones [37]. Here we follow an approximate but simpler approach for the case $L_{\text{x}}=L_{\text{y}}=L_{\text{z}}=L$. The modes in (25) contributing to the EM field are those with index combinations $(n_{x},n_{y},n_{z})$, corresponding to the wavenumber points ${\bm{\kappa}}_{n_{x},n_{y},n_{z}}=(k_{n_{x}},k_{n_{y}},k_{n_{z}})$, for which the intersection with the surface of radius $k_{0}$ defined in (6) is not negligible. Their number can be approximatively evaluated as the number of wavenumber points ${\bm{\kappa}}_{n_{x},n_{y},n_{z}}$ falling within the shell of radius $\kappa_{0}$ and thickness $\delta$ (see Fig. 6). Denote with ${\cal{P}}_{3}$ the set of indexes $(n_{x},n_{y},n_{z})$ whose wavenumber satisfies this condition. For instance, $\delta$ can be taken equal to the first lobe of the Sinc function, i.e., $\delta=2\pi/L$. Specifically, it is $DoF\simeq|{\cal{P}}_{3}|=\nu\cdot V_{\delta}$, where $\nu=\left(\frac{L}{2\pi}\right)^{3}$ is the density of the wavenumber points and $V_{\delta}$ is the volume of the shell. Therefore, it is
$DoF\simeq\nu\cdot V_{\delta}=\frac{\pi}{3}+\frac{4\pi L^{2}}{\lambda^{2}}\simeq\frac{\pi^{2}L^{2}}{\lambda^{2}}$ (26)
for $L\gg\lambda$, which is the same result one can obtain using a more precise but complex approach based on spherical harmonic waves. Note that using volumetric antennas than planar antennas a DoF gain of $\pi$ can only be obtained. It can be pointed out that while for the 1D case sampling the current with spacing $\lambda/2$ is necessary to represent the current components contributing to the radiating EM field, in the 2D and 3D cases sampling at $\lambda/2$ is only a sufficient but no longer necessary condition. Readers can refer to [41] for a more detailed discussion.
@@ IV-B DoF of the Single Link
Consider a communication system involving a transmitting space $\mathcal{S}_{\text{T}}$ and a receiving space $\mathcal{S}_{\text{R}}$. Each space allows for a maximum number of DoF in the unbounded space, as discussed in Sec. IV-A. However, only a subset of these available DoF can be exploited for communication, corresponding to the $N$ well-coupled communication modes with eigenvalues $\sigma_{n}^{2}$ described in Sec. III-B. When currents are excited at $\mathcal{S}_{\text{T}}$ corresponding to these well-coupled modes, the resulting EM field is predominantly concentrated on the receiver space $\mathcal{S}_{\text{R}}$. In contrast, exciting modes outside this subset of well-coupled modes leads to an EM field largely dispersed away from $\mathcal{S}_{\text{R}}$.
@@
In the context of communication systems, the $N$ communication modes represent the available DoF between the transmitting space $\mathcal{S}_{\text{T}}$ and the receiving space $\mathcal{S}_{\text{R}}$, which can be modeled using the communication scheme shown in Fig. 7. Here, $N$ information symbols ${\bf{x}}\in\mathbb{C}^{N}$ are associated with the basis functions $\left\{{\bm{\varphi}}_{n}({\bf{r}})\right\}$, forming the source current density ${\mathsf{J}}({\bf{r}})=\sum_{n}x_{n}{\bm{\varphi}}_{n}({\bf{r}})$ on $\mathcal{S}_{\text{T}}$. At the receiver, the information is recovered by correlating the received EM field ${\mathsf{E}}({\bf{r}})$ with the corresponding basis functions $\left\{{\mathbf{\Psi}}_{n}({\bf{r}})\right\}$ on $\mathcal{S}_{\text{R}}$, i.e., $y_{n}=\left<{{\mathsf{E}}({\bf{r}})}\,,{{\mathbf{\Psi}}_{n}({\bf{r}})}\right>$, resulting in a set of matched filters operating in the spatial domain. Formally, considering the effect of noise, the received signal $\mathbf{y}\in\mathbb{C}^{N}$ is given by the following equation
$\mathbf{y}=\bm{\Sigma}\,\mathbf{x}+\mathbf{n}$ (27)
@@
$C=\mathop{\max}_{\begin{subarray}{c}p_{1}\geq 0,\ldots,p_{N}\geq 0\\ \mathrm{tr}(\mathbf{Q})\leq P_{\text{T}}\end{subarray}}\sum_{n=1}^{N}\log_{2}\left(1+\frac{p_{n}\sigma_{n}^{2}}{\sigma_{\mathrm{noise}}^{2}}\right)$ (28)
@@
$p_{n}=\max\left(\mu-\frac{\sigma_{\mathrm{noise}}^{2}}{\sigma_{n}^{2}},0\right),\quad\quad n=1,\ldots,N$ (29)
@@ IV-B1 DoF for 1D Structures
Consider two 1D segments, as described in Sec. IV-A, placed parallel to each other and aligned along their centers (i.e., in a paraxial configuration). We focus on a single polarization. The lengths of the segments are denoted as $L_{\mathrm{T}}$ and $L_{\mathrm{R}}$, and the distance between their centers is $d$. It can be demonstrated that the $N$ most significant eigenvalues in this configuration are nearly identical in magnitude and then rapidly decay. Therefore, according to the water-filling principle in (29), the power can be evenly distributed among the $N$ modes. The number of DoF of the link, or the number of significant eigenvalues, can thus be determined for a single polarization as follows [39]
@@
This classical expression, while accurately determining the number of DoF when $L_{\mathrm{T}},L_{\mathrm{R}}\gg d$ and the sizes of $L_{\mathrm{T}}$ and $L_{\mathrm{R}}$ are nearly comparable, becomes inadequate when one of the segments grows excessively large. In fact, (30) appears to suggest that the DoF can increase indefinitely if either $L_{\mathrm{T}}$ or $L_{\mathrm{R}}$ becomes infinitely large. However, this contradicts the results obtained in Sec. IV-A, where the maximum number of DoF is bounded and cannot exceed the values derived for an unbounded space. A more accurate expression was derived in [43] by analyzing the scalar Green operator and modeling the communication modes as focused beams, which leads to
$N=\frac{2L_{\mathrm{T}}}{\lambda}\zeta\,$ (31)
with $\zeta=L_{\mathrm{R}}/\sqrt{4d^{2}+L_{\mathrm{R}}^{2}}$. From expression (31), it can be observed that as $L_{\mathrm{R}}\to\infty$, we have $\zeta\to 1$, and thus the limit is $N=\frac{2L_{\mathrm{T}}}{\lambda}$, which corresponds to the actual limit in the unbounded space presented in Sec. IV-A. Among the $N=\frac{2L_{\mathrm{T}}}{\lambda}$ orthogonal beamsteering directions associated with (23) that can be generated with the segment of length $L_{\mathrm{T}}$, (31) indicates the number of beams that intersect the parallel segment of length $L_{\mathrm{R}}$ at a distance $d$.
@@ IV-B2 DoF for 2D Structures
Consider now two square surfaces with areas, respectively, $A_{\mathrm{T}}=L_{\mathrm{T}}^{2}$ and $A_{\mathrm{R}}=L_{\mathrm{R}}^{2}$, where $A_{\mathrm{T}}<A_{\mathrm{R}}$. These surfaces are located at a distance $d$ in a paraxial configuration and outside the reactive region of the transmitting area. When $d>L_{\mathrm{R}}$, the classical expression for the available DoF (per polarization) is [39]
$N=\frac{A_{\mathrm{T}}A_{\mathrm{R}}}{\lambda^{2}d^{2}}\,.$ (32)
In [5], the following alternative expression is derived for $A_{\mathrm{T}}<A_{\mathrm{R}}$
@@
Unlike (32), (33) provides the actual number of DoF for squared planar paraxial surfaces of any size, also demonstrating a saturation when $d\ll L_{\mathrm{R}}$. In the case $L_{\mathrm{R}}\rightarrow\infty$, the limit is $N=\frac{\pi A_{\mathrm{T}}}{\lambda^{2}}$, which matches the result obtained for the unbounded space in Sec. IV-A.
IV-C DoF of the Cascade Link
When a reconfigurable scattering device is placed between the transmitter and the receiver, as discussed in Sec. III-C and shown in Fig. 3, we define $N$ and $M$ as the maximum number of “usable” channels (i.e., communication modes) into and out of the scattering device, respectively. Here, $N$ and $M$ correspond to the rank of $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$, respectively, with $N,M\leq K$. The maximum number of end-to-end communication modes is given by $R=\min(N,M)$. The equivalent end-to-end channel can be written as
$\mathbf{y}=\mathbf{H}(\bm{\theta})\,\mathbf{x}+\mathbf{n}$ (34)
where now ${\bf{y}},{\bf{x}},{\bf{n}}\in\mathbb{C}^{K}$, and the end-to-end channel matrix ${\bf{H}}(\theta)$ defined in (17) can be written equivalently as ${\bf{H}}(\theta)=\mathbf{G}_{\text{R}}{\bf{U}}{\bf{G}}^{-1}{\bf{R}}(\bm{\theta}){\bf{V}}^{-1}\mathbf{G}_{\text{T}}={\bf{H}}_{\mathrm{R}}{\bf{R}}(\bm{\theta}){\bf{H}}_{\mathrm{T}}$, with ${\bf{H}}_{\mathrm{R}}=\mathbf{G}_{\text{R}}{\bf{U}}{\bf{G}}^{-1}$ and ${\bf{H}}_{\mathrm{T}}={\bf{V}}^{-1}\mathbf{G}_{\text{T}}$, to explicit the effect of the reflection matrix ${\bf{R}}(\bm{\theta})$.
It is of interest to investigate the optimal configuration of the reconfigurable scattering device through the matrix ${\bf{D}}(\bm{\theta})$ (thus, ${\bf{R}}(\bm{\theta})$ and $\mathbf{H}(\bm{\theta})$) maximizing the end-to-end channel capacity given $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$ under some constraint on the matrix ${\bf{R}}(\bm{\theta})\approx{\bf{G}}\,{\bf{D}}(\bm{\theta})$ (we make use of the Born approximation), for example, a lossless constraint for the reconfigurable scattering device [^5] In this case, the power of the incident electric field ${\bf{e}}_{\mathrm{i}}$ equals that of the scattered electric field ${\bf{e}}_{\mathrm{s}}={\bf{R}}(\bm{\theta})\,{\bf{e}}_{\mathrm{i}}$. leading to ${{\bf{R}}(\bm{\theta})\,{\bf{R}}^{{\mathrm{H}}}(\bm{\theta})={\bf{I}}_{K}}$.

### v3

@@ IV-A2 DoF for 2D Structures
In this second case, $\cal{S}$} is a surface of size $L_{\text{x}}\times L_{\text{y}}$. Similarly to the 1D case, the current density can be represented as a 2D series expansion of Fourier basis functions. In the wavenumber domain, we have [15]
${\widetilde{J}}(\kappa_{x},\kappa_{y})=\sum_{n_{x},n_{y}}j_{n_{x},n_{y}}\,\sqrt{L_{\text{x}}\,L_{\text{y}}}\,\text{sinc}\left({\frac{L_{\text{x}}}{2\pi}\left(\kappa_{x}-k_{n_{x}}\right)}\right)\text{sinc}\left({\frac{L_{\text{y}}}{2\pi}\left(\kappa_{y}-k_{n_{y}}\right)}\right)\,$, (25)
where $k_{n_{x}}=2\,n_{x}\,\pi/L_{\text{x}}$, and $k_{n_{y}}=2\,n_{y}\,\pi/L_{\text{y}}$. Note that only the set of indexes $(n_{x},n_{y})\in{\cal{P}}_{2}$, with ${\cal{P}}_{2}=\left\{(n_{x},n_{y}):\left(\frac{n_{x}\lambda}{L_{\text{x}}}\right)^{2}+\left(\frac{n_{y}\lambda}{L_{\text{y}}}\right)^{2}\leq 1\right\}$, satisfying $k_{n_{x}}^{2}+k_{n_{y}}^{2}\leq k_{0}^{2}$, are not evanescent, i.e., lie in the visible region. For instance, for a square surface of size $L_{\text{x}}=L_{\text{y}}=L$, it is ${N_{\text{u}}}\simeq|{\cal{P}}_{2}|\simeq\frac{\pi L^{2}}{\lambda^{2}}$ (per polarization) [15]. Interestingly, this is $\pi/4=0.79$ of the square of the DoF of a linear segment. As a consequence, by drawing a parallel as before with uniform planar arrays, we have that a ${N_{\text{a}}}\times{N_{\text{a}}}$ planar array will not be able to generate ${N_{\text{a}}}^{2}$ orthogonal beams (i.e., $\frac{4L^{2}}{\lambda^{2}}$), as could be erroneously suggested by intuition, but only a subset ${N_{\text{u}}}<{N_{\text{a}}}^{2}$, specifically ${N_{\text{u}}}=\frac{\pi}{4}{N_{\text{a}}}^{2}\approx 0.79\,{N_{\text{a}}}^{2}$ beams satisfying the condition of non-evanescent modes.
@@ IV-A3 DoF for 3D Structures
${\widetilde{J}}(\kappa_{x},\kappa_{y},\kappa_{z})=\hskip-8.5359pt\sum_{n_{x},n_{y},n_{z}}\hskip-8.5359ptj_{n_{x},n_{y},n_{z}}\sqrt{L_{\text{x}}L_{\text{y}}L_{\text{z}}}\,\text{sinc}\left({\frac{L_{\text{x}}}{2\pi}\left(\kappa_{x}-k_{n_{x}}\right)}\right)\text{sinc}\left({\frac{L_{\text{y}}}{2\pi}\left(\kappa_{y}-k_{n_{y}}\right)}\right)\text{sinc}\left({\frac{L_{\text{z}}}{2\pi}\left(\kappa_{z}-k_{n_{z}}\right)}\right)\,$, (26)
@@
[…]
The derivation of the number of DoF for a 3D structure is more involved. The typical approach is to consider the smallest sphere enclosing the space $\cal{S}$}, define a complete basis set given by the spherical harmonic functions on its surface, and identify the $N$ dominant ones [35]. Here we follow an approximate but simpler approach for the case $L_{\text{x}}=L_{\text{y}}=L_{\text{z}}=L$. The modes in (26) contributing to the EM field are those with index combinations $(n_{x},n_{y},n_{z})$, corresponding to the wavenumber points ${\bm{\kappa}}_{n_{x},n_{y},n_{z}}=(k_{n_{x}},k_{n_{y}},k_{n_{z}})$, for which the intersection with the surface of radius $k_{0}$ defined in (7) is not negligible. Their number can be approximatively evaluated as the number of wavenumber points ${\bm{\kappa}}_{n_{x},n_{y},n_{z}}$ falling within the shell of radius $\kappa_{0}$ and thickness $\delta$ (see Fig. 5). Denote with ${\cal{P}}_{3}$ the set of indexes $(n_{x},n_{y},n_{z})$ whose wavenumber satisfies this condition. For instance, $\delta$ can be taken equal to the first lobe of the sinc function, i.e., $\delta=2\pi/L$. Specifically, it is ${N_{\text{u}}}\simeq|{\cal{P}}_{3}|=\nu\cdot V_{\delta}$, where $\nu=\left(\frac{L}{2\pi}\right)^{3}$ is the density of the wavenumber points and $V_{\delta}$ is the volume of the shell. Therefore, it is
${N_{\text{u}}}\simeq\nu\cdot V_{\delta}=\frac{\pi}{3}+\frac{4\pi L^{2}}{\lambda^{2}}\simeq\frac{\pi^{2}L^{2}}{\lambda^{2}}\,$, (27)
for $L\gg\lambda$, which is the same result one can obtain using a more precise but complex approach based on spherical harmonic waves. Note that, when using volumetric antennas instead of planar antennas, a DoF gain of at most $\pi$ can be achieved. It can be pointed out that while for the 1D case sampling the current with spacing $\lambda/2$ is necessary to represent the current components contributing to the radiating EM field, in the 2D and 3D cases sampling at $\lambda/2$ is only a sufficient but no longer necessary condition. Readers can refer to [40] for a more detailed discussion.
@@ IV-B DoF of the Single Link
Consider a communication system involving a transmitting space $\mathcal{S}_{\text{T}}$ and a receiving space $\mathcal{S}_{\text{R}}$. Each space allows for a maximum number of DoF in the unbounded space, as discussed in Section IV-A. However, only a subset of these available DoF can be exploited for communication, corresponding to the $N$ well-coupled communication modes with eigenvalues $\sigma_{n}^{2}$ described in Section III-B. When currents are excited at $\mathcal{S}_{\text{T}}$ corresponding to these well-coupled modes, the resulting EM field is predominantly concentrated on the receiver space $\mathcal{S}_{\text{R}}$. In contrast, exciting modes outside this subset of well-coupled modes leads to an EM field largely dispersed away from $\mathcal{S}_{\text{R}}$.
@@
In the context of communication systems, the $N$ communication modes represent the available DoF between the transmitting space $\mathcal{S}_{\text{T}}$ and the receiving space $\mathcal{S}_{\text{R}}$, which can be modeled using the communication scheme shown in Fig. 6. Here, $N$ information symbols ${\mathbf{x}}\in\mathbb{C}^{N}$ are associated with the basis functions $\left\{{\bm{\varphi}}_{n}({\mathbf{r}})\right\}$, forming the source current density ${\mathsf{J}}({\mathbf{r}})=\sum_{n}x_{n}{\bm{\varphi}}_{n}({\mathbf{r}})$ on $\mathcal{S}_{\text{T}}$. At the receiver, the information is recovered by correlating the received EM field ${\mathsf{E}}({\mathbf{r}})$ with the corresponding basis functions $\left\{{\mathbf{\Psi}}_{n}({\mathbf{r}})\right\}$ on $\mathcal{S}_{\text{R}}$, i.e., $y_{n}=\left<{{\mathsf{E}}({\mathbf{r}})}\,,{{\mathbf{\Psi}}_{n}({\mathbf{r}})}\right$>, resulting in a set of matched filters operating in the spatial domain. Formally, considering the effect of noise, the received signal $\mathbf{y}\in\mathbb{C}^{N}$ is given by the following equation
$\mathbf{y}=\bm{\Sigma}\,\mathbf{x}+\mathbf{n}\,$, (28)
@@
$C=\mathop{\max}_{\begin{subarray}{c}p_{1}\geq 0,\ldots,p_{N}\geq 0\\ \mathrm{tr}(\mathbf{Q})\leq P_{\text{T}}\end{subarray}}\sum_{n=1}^{N}\log_{2}\left(1+\frac{p_{n}\sigma_{n}^{2}}{\sigma_{\mathrm{noise}}^{2}}\right)\,$, (29)
@@
$p_{n}=\max\left(\mu-\frac{\sigma_{\mathrm{noise}}^{2}}{\sigma_{n}^{2}},0\right),\quad\quad n=1,\ldots,N\,$, (30)
@@ IV-B1 DoF for 1D Structures
Consider two 1D segments as described in Section IV-A, placed parallel to each other and aligned along their centers (i.e., in a paraxial configuration). We focus on a single polarization. The lengths of the segments are denoted as $L_{\mathrm{T}}$ and $L_{\mathrm{R}}$, and the distance between their centers is $d$. It can be demonstrated that the $N$ most significant eigenvalues in this configuration are nearly identical in magnitude and then rapidly decay. Therefore, according to the water-filling principle in (30), the power can be evenly distributed among the $N$ modes. The number of DoF of the link, or the number of significant eigenvalues, can thus be determined for a single polarization as follows [38]
@@
This classical expression, while accurately determining the number of DoF when $L_{\mathrm{T}},L_{\mathrm{R}}\gg d$ and the sizes of $L_{\mathrm{T}}$ and $L_{\mathrm{R}}$ are nearly comparable, becomes inadequate when one of the segments grows excessively large. In fact, (31) appears to suggest that the DoF can increase indefinitely if either $L_{\mathrm{T}}$ or $L_{\mathrm{R}}$ becomes infinitely large. However, this contradicts the results obtained in Section IV-A, where the maximum number of DoF is bounded and cannot exceed the values derived for an unbounded space. A more accurate expression was derived in [42] by analyzing the scalar Green’s operator and modeling the communication modes as focused beams, which leads to
$N=\frac{2L_{\mathrm{T}}}{\lambda}\zeta\,$, (32)
with $\zeta=L_{\mathrm{R}}/\sqrt{4d^{2}+L_{\mathrm{R}}^{2}}$. From expression (32), it can be observed that as $L_{\mathrm{R}}\to\infty$, we have $\zeta\to 1$, and thus the limit is $N={N_{\text{u}}}=\frac{2L_{\mathrm{T}}}{\lambda}$ which corresponds to the actual limit in the unbounded space presented in Section IV-A. Among the ${N_{\text{u}}}=\frac{2L_{\mathrm{T}}}{\lambda}$ orthogonal beamsteering directions associated with (24) that can be generated with the segment of length $L_{\mathrm{T}}$, (32) indicates the number of beams that intersect the parallel segment of length $L_{\mathrm{R}}$ at a distance $d$.
@@ IV-B2 DoF for 2D Structures
Consider now two square surfaces with areas, respectively, $A_{\mathrm{T}}=L_{\mathrm{T}}^{2}$ and $A_{\mathrm{R}}=L_{\mathrm{R}}^{2}$. These surfaces are located at a distance $d$ in a paraxial configuration and outside the reactive region of the transmitting area. When $d\gg L_{\mathrm{T}},L_{\mathrm{R}}$, the classical expression for the available DoF (per polarization) is [38]
$N=\frac{A_{\mathrm{T}}A_{\mathrm{R}}}{\lambda^{2}d^{2}}={N_{\text{u}}}^{(\text{T})}\,{N_{\text{u}}}^{(\text{R})}\,\left(\frac{\lambda}{\pi d}\right)^{2}\leq\min\left({N_{\text{u}}}^{(\text{T})},{N_{\text{u}}}^{(\text{R})}\right)$ (33)
where ${N_{\text{u}}}^{(\text{T})}$ and ${N_{\text{u}}}^{(\text{R})}$ are the number of DoF in the unbounded space of the transmitting and receiving areas, respectively. In [22], the following alternative expression is derived for $A_{\mathrm{T}}<A_{\mathrm{R}}$
@@
Unlike (33), (34) provides the actual number of DoF for squared planar paraxial surfaces of any size, also demonstrating a saturation when $d\ll L_{\mathrm{R}}$. In the case $L_{\mathrm{R}}\rightarrow\infty$, the limit is $N={N_{\text{u}}}^{(\text{T})}=\frac{\pi A_{\mathrm{T}}}{\lambda^{2}}$, which matches the result obtained for the unbounded space in Section IV-A.
IV-C DoF of the Cascade Link including a Reconfigurable Scattering Device
When a reconfigurable scattering device is placed between the transmitter and the receiver, as discussed in Section III-C and shown in Fig. 3, we define $N$ and $M$ as the maximum number of “usable” channels (i.e., communication modes) into and out of the scattering device, respectively. Here, $N$ and $M$ correspond to the rank of $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$, respectively, with $N,M\leq K$. The maximum number of end-to-end communication modes is given by ${N_{\text{c}}}=\min(N,M)$. Note that, since $N,M\leq{N_{\text{u}}}$, where $N_{\text{u}}$} is the number of DoF of the reconfigurable scattering device in the unbounded space derived in Section IV-A, it is ${N_{\text{c}}}\leq{N_{\text{u}}}$. The equivalent end-to-end channel can be written as
$\mathbf{y}=\mathbf{H}(\bm{\theta})\,\mathbf{x}+\mathbf{n}\,$, (35)
where now ${\mathbf{y}},{\mathbf{x}},{\mathbf{n}}\in\mathbb{C}^{K}$, and the end-to-end channel matrix ${\mathbf{H}}(\theta)$ defined in (18) can be written equivalently as ${\mathbf{H}}(\theta)=\mathbf{G}_{\text{R}}{\mathbf{U}}{\mathbf{G}}^{-1}{\mathbf{R}}(\bm{\theta}){\mathbf{V}}^{-1}\mathbf{G}_{\text{T}}={\mathbf{H}}_{\mathrm{R}}{\mathbf{R}}(\bm{\theta}){\mathbf{H}}_{\mathrm{T}}$, with ${\mathbf{H}}_{\mathrm{R}}=\mathbf{G}_{\text{R}}{\mathbf{U}}{\mathbf{G}}^{-1}$ and ${\mathbf{H}}_{\mathrm{T}}={\mathbf{V}}^{-1}\mathbf{G}_{\text{T}}$, to explicit the effect of the reflection matrix ${\mathbf{R}}(\bm{\theta})$.
It is of interest to investigate the optimal configuration of the reconfigurable scattering device through the matrix ${\mathbf{D}}(\bm{\theta})$ (thus, ${\mathbf{R}}(\bm{\theta})$ and $\mathbf{H}(\bm{\theta})$) maximizing the end-to-end channel capacity given $\mathbf{G}_{\text{T}}$ and $\mathbf{G}_{\text{R}}$ under some constraint on the matrix ${\mathbf{R}}(\bm{\theta})\approx{\mathbf{G}}\,{\mathbf{D}}(\bm{\theta})$ (we make use of the Born approximation), for example, a lossless constraint for the reconfigurable scattering device [^6] In this case, the power of the incident electric field ${\mathbf{e}}_{\mathrm{i}}$ equals that of the scattered electric field ${\mathbf{e}}_{\mathrm{s}}={\mathbf{R}}(\bm{\theta})\,{\mathbf{e}}_{\mathrm{i}}$. leading to ${\mathbf{R}}(\bm{\theta})\,{\mathbf{R}}^{{\mathrm{H}}}(\bm{\theta})={\mathbf{I}}_{K}$}.